what is the time required to pass 1 F of electricity through electrolyte with a current of 0.35 ampere


1
Expert's answer
2022-02-02T04:28:05-0500

Solution:

A Faraday is defined as amount of electric charge in one mole of electrons.

1 F = 96485 coulombs


The amount of charge passed through electrolyte can be calculated using the equation:

charge = current × time

Q = I × t

where,

  • charge (Q) is measured in coulombs (C)
  • current (I) is measured in ampers (A)
  • time (t) is measured in seconds (s)

Thus:

t = Q / I

t = (96485 C) / (0.35 A) = 275671.43 s

275671.43 s = 76 hr 34 min 31 sec = 76.58 hrs


Answer: 76.58 hrs